In Discussion #1, you calculated the volume of gas that your explosive would produce if 1 g of it was used, assuming standard temperature and pressure (STP)and any water produced in the gaseous state. Is this actually accurate, as the gas is usually heated as a result of the explosion, resulting in a much larger volume of gas? But what temperature will the gas reach? Calculate the final volume of the gas produced from 1.00 g of your explosive using Charles' Law, and assuming your explosion reaches a temperature of 3000 K.  

Remember Charles' law? V1/T1 = V2/T2

Explosive is TATB and volume of gas from previous answer =0.78L
Answer #1

Solution :-

Initial volume V1 = 0.78 L

Initial temperature T1 = 273 K ( standard temperature)

Final temperature T2 = 3000 K

Final volume V2 = ?

Charles law formula

V1/T1 = V2/T2

V2 = V1*T2/T1

V2 = 0.78 L * 3000 K / 273 K

V2 = 8.57 L

Therefore the volume of the gases would be 8.57 L